OpenStudy (anonymous):

Assume angle C is a right angle and given the conditions b=2 and A= 30 degrees solve for the side lengths of the triangle.

OpenStudy (anonymous):

So we know this is a 30,60,90 triangle which has special properties yes?

OpenStudy (anonymous):

@Faith_Rochelle

OpenStudy (anonymous):

The easiest way to do this is to just draw it out is there a picture with the question?

OpenStudy (anonymous):

no there isn't.

OpenStudy (anonymous):

ok, so I would draw this out which will make it much simpler

OpenStudy (anonymous):

can you draw this triangle based on the info above?

OpenStudy (anonymous):

i think traditionally b is the line in between 90 and 30 so it should be the line between angles A and C

OpenStudy (anonymous):

I don't understand where to draw it at.

OpenStudy (anonymous):

so the special rules for 30,60,90 are: a = a b = a√3 c= 2a

OpenStudy (anonymous):

ok so you draw a 30,60,90 triangle the line connecting the 90 degree angle and the 30 degree angle is b, 2

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):



OpenStudy (anonymous):

the bottom right angle should be 60 degrees

OpenStudy (anonymous):

and you have to switch a and b

OpenStudy (anonymous):

then refer to that for values of sides should be pretty straight forward happy to help though :)

OpenStudy (anonymous):

Is this making sense to you? we know triangle is a 30 60 90 because they tell us one side is 90 one is 30, which means the other has to be 60 to add to 180 and make a triangle. So these triangles have special rules about the length of their sides and how they compare to each other I posted a picture of the above. So we kno the b value or a√3 is equal to 2

OpenStudy (anonymous):

so we make an equation b = a√3 and plug our value in 2 = a√3 and solve for the a side. Are you following me this far?

OpenStudy (anonymous):

No. What do we plug in for A?

OpenStudy (anonymous):

ok you left but solve for a and then plug that into c which is c = 2a and you will have values for a,b, and c

OpenStudy (anonymous):

we do not know a we solve it with the eq i explained above...

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

I don't see an equation for a..

OpenStudy (anonymous):

there isnt an equation for a it is the value of one side and the value of the other are based off of a

OpenStudy (anonymous):

but we work backwords because we know b to find a, and can use that to find c

OpenStudy (anonymous):

ok here, b = 2 a = ? c = ? b = a√3 so, 2 = a√3 a = 2/√3 c = 2a so, c = 2(2/√3) c = 4/√3 a = 2/√3 b = 2 c = 4/√3

OpenStudy (anonymous):

I would ask your teacher to help explain i dont know what else i can do through the web good luck faith hope you get it :)